OpenStudy (anonymous):

5 (x • y) = (5 • x) y is an example of which algebraic property?

OpenStudy (ghazi):

associative

OpenStudy (anonymous):

\[a \cdot (b \cdot c) = (a \cdot b) \cdot c\] This property is called Associative Property of Multiplication..